OpenStudy (anonymous):

the third floor of a house is 8m above street level. How much work is needed to move a 150kg refrigerator to the third floor

OpenStudy (anonymous):

Work if a measure of energy. The work needed actually equals the gain in (gravitational) potential energy by the refrigerator (if we assume both the initial and final velocities of the refrigerator are 0). So Work=Potential Energy Gained=m*g*h=150*9.8*8=1,18*10^4 J

OpenStudy (shamim):

i think the result will b 11760 joule

OpenStudy (shamim):

did u get this solution@ chelsie

OpenStudy (shamim):

we know work done w =mgh

OpenStudy (anonymous):

it's either -11760 or zero, the work will be negative because you're acting against gravity or it will be zero, because by work definition, it is w= f*d* cos theta. since our angle is 90, cos of 90 is zero.

OpenStudy (shamim):

no

OpenStudy (shamim):

sorry i m not agree with u

OpenStudy (shamim):

i think work done by an external force is 11760 joule. this work done is positive. because the angle between external force and displacement is 0 degree. and cos0=1. we know that work done \[W=F \times s\]

OpenStudy (shamim):

\[or, w=FScos \theta\]

OpenStudy (shamim):

\[or, W=FS \cos0=FS \times 1\]

OpenStudy (shamim):

\[or, W=FS\]

OpenStudy (shamim):

\[or, W=mgS\]

OpenStudy (shamim):

\[or, W=150 \times 9.8 \times 8=11760 joule\]

OpenStudy (shamim):

but if we wanna calculate the work done by gravitational force

OpenStudy (shamim):

\[W=FScos \theta= FS \cos 180=FS(-1)=-FS=-mgs=-150 \times 9.8 \times 8\]

OpenStudy (shamim):

\[or, W=-11760 joule\]

OpenStudy (shamim):

\[\theta =180\]

OpenStudy (shamim):

because the gravitational force is directed along downward and displacement is upward. so the angle between gravitational force and displacement is 180 degree

OpenStudy (anonymous):

If you're talking about the work done by gravity, then the result is negative. But if you're talking about the work done by an elevator, then the work if obviously positive (think about it, you do need put a lot of effort to lift such a mass).

OpenStudy (anonymous):

And the answer of 11760J is indeed true, if you consider that you can somehow lift the mass without friction and in a straight vertical line.